OpenStudy (anonymous):

Can someone please help me with this? What is the volume of the solid generated by revolving the region bounded by the given lines and curves about the x-axis? y= sq root(3x) y=3, x=0

OpenStudy (anonymous):

maybe x=3 not y=3

OpenStudy (anonymous):

no it is y=3

OpenStudy (anonymous):

\[\int\limits_{0}^{3}3*\pi*xdx=1.5x^2*\pi|=1.5*9*\pi\] i dont know how to solve if there y=3

OpenStudy (dumbcow):

 \[V = \pi \int\limits_{0}^{3}R^{2} -r^{2} dx\] \[V = \pi \int\limits_{0}^{3}3^{2} -(\sqrt{3x})^{2} dx\]